OpenStudy (anonymous):

i missed my class... how do i solve this? Square root of 110 - n(110 and - n are all under the sqr root) =n so Sqr of 110 -n = n

OpenStudy (anonymous):

sqrt(110-n)=n 110-n=n² n²+n-110=0 (n+11)(n-10)=0 n+11=0 or n-10 = 0 n=-11 or n=10

OpenStudy (anonymous):

\[\sqrt{110 - n} = n\]